Your uncle is considering investing in a new company that will produce high quality stereo speakers. The sales price would be set at it; the variable cost per unit is estimated to be $60.00; and fixed costs are estimated at $1,000,000. 1.7 times the variable cost per un What sales volume would be required to break even, .e., to have EBIT - zero? O A. 33333.33 O B. 16666.67 o C. 15697.43 O D.28333.33 O E. 23809.52
Answer #1

Variable cost per unit=60

Sales price per unit=1.7*60=102

Breakeven units=1000000/(102-60)=23809.52

Option E